# Marginally interpretable spatial logistic regression with bridge processes

### v2

@@ 1 Introduction
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid x_{ij},u(s_{i})\}]=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i}),\quad u(\cdot)\sim\textrm{mean zero process},$ (1)
@@
where $Y_{ij}$ is the $j$th binary observation at the $i$th spatial location $s_{i}$ ($i=1,\dots,n$, $j=1,\dots,n_{i}$), $x_{ij}\in\mathbb{R}^{p}$ are covariates, and $\{u(s):s\in\mathscr{S}\}$ is a mean zero stochastic process defined on a spatial domain $\mathscr{S}$, with the most common choice being the Gaussian process. This generalized linear mixed effects model formulation (Diggle et al.,, 1998) provides a full characterization of the data, naturally accounting for spatial dependence, and allowing for prediction at new locations in a coherent probabilistic framework.
@@
When interpreting the regression coefficient $\beta$, the logit link provides a familiar understanding in terms of log odds ratios. However, the inclusion of random effects changes the meaning of $\beta$ from a marginal effect on the population to a conditional effect specific to the site. Specifically, when the normal random effect $u$ is integrated out from (1), the induced model for $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}$ is no longer a linear function of $x_{ij}$ (Zeger et al.,, 1988). There exists a rich literature on methods designed for marginally specified models, most notably generalized estimating equations (Liang and Zeger,, 1986), but such approaches cannot provide probabilistic prediction at new locations. This creates difficulty for researchers who wish to maintain a population-averaged interpretation of $\beta$ while also accounting for spatial dependence and making predictions. We refer the reader to Neuhaus et al., (1991); Heagerty and Zeger, (2000); Hubbard et al., (2010) for a detailed comparison of marginal and conditional models for clustered binary data.
@@
For logistic models with independent random intercepts, Wang and Louis, (2003) proposed a family of univariate bridge distributions as an alternative to normal random effects. In marginalizing random intercept logistic regression models over the bridge distribution, the resulting model also has a logistic form so that $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}$ is a linear function of $x_{ij}$ with coefficients proportional to $\beta$. Thus, bridge-distributed random effects allow both a conditional and marginal interpretation of regression coefficients, motivating applications in many different contexts (Bandyopadhyay et al.,, 2010; Tu et al.,, 2011; Asar,, 2021). However, the bridge distribution has not been naturally extended to multivariate settings for correlated random effects.
@@
We propose a new class of marginally interpretable spatial logistic regression models based on a novel spatial bridge process. We identify the normal-scale mixture representation of the bridge distribution and propose its multivariate extension. In contrast to existing copula-based constructions that have bridge-distributed marginals (Lin et al.,, 2010; Li et al.,, 2011; Parzen et al.,, 2011; Boehm et al.,, 2013; Swihart et al.,, 2014), the bridge process has appealing properties, such as transparent correlation structure and comes with significant computational benefits.
@@ 2 Marginally interpretable spatial logistic regression models
2.1 Bridge distribution as a normal scale mixture
The bridge distribution ${p_{\textsc{br}}}(u;\phi)$ with parameter $\phi\in(0,1)$ is derived from the solution of the integral identity $\int_{-\infty}^{\infty}\mathrm{logit}^{-1}(\eta+u){p_{\textsc{br}}}(u;\phi)\mathrm{d}u=\mathrm{logit}^{-1}(\phi\eta)$, see Figure 1 (left) for a comparison with the normal distribution. This ensures that the induced marginal model, when random intercepts are integrated out, is also a logistic regression model with coefficients scaled by $\phi$.
We aim to design mean-zero stochastic processes $\{u(s):s\in\mathscr{S}\}$ for random effects, providing a dual interpretation of coefficients with appealing properties. This involves developing a new process whose finite-dimensional realizations have bridge-distributed marginals. We first identify a scale mixture of normal representation of the bridge distribution, which forms the basis of our construction for multivariate extensions. We defer all proofs to Appendix A.1.
@@ 2.1 Bridge distribution as a normal scale mixture
with the mixing variable $\lambda$ with density ${p_{\textsc{m}}}(\lambda;\phi)$ is equal in distribution to $2\phi^{-2}\sum_{k=1}^{\infty}A_{k}B_{k}/k^{2}$, where $A_{k}\sim\mathrm{Exp}(1)$, $B_{k}\sim\mathrm{Ber}(1-\phi^{2})$, independently for $k\in\mathbb{N}$.
The proof is similar to West, (1987). The mixing distribution has a similar form to the squared Kolmogorov distribution that serves as a normal variance mixing distribution of the logistic distribution (Andrews and Mallows,, 1974), except for the presence of Bernoulli random variables $\{B_{k}\}$. More details of the mixing distribution can be found in Appendix A.3.
@@ 2.2 Bridge processes for logit link
We now introduce a process and multivariate extension of the univariate bridge distribution, which is uniquely determined by $\phi$ and the choice of a correlation kernel.
Definition 1.
Let $R$ be an $n\times n$ positive semidefinite correlation matrix. We say $u$ follows an $n$-dimensional multivariate bridge distribution for logit link with parameters $\phi\in(0,1)$ and $R$ if
$u\,|\,\lambda\sim\mathrm{N}_{n}\left(0,\lambda R\right),\quad\lambda\sim{p_{\textsc{m}}}(\phi).$ (3)
Furthermore, let $\mathcal{R}:\mathscr{S}\times\mathscr{S}\to[-1,1]$ be a positive semidefinite kernel with $\mathcal{R}(s,s)=1$ for every $s\in\mathscr{S}$. We say $\{u(s)\in\mathbb{R}:s\in\mathscr{S}\}$ is a bridge process with parameter $\phi$ and correlation kernel $\mathcal{R}$ if every finite collection $\{u(s_{1}),\dots,u(s_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows an $n$-dimensional multivariate bridge distribution with parameters $\phi$ and $R$ with $(i,j)$th element $R_{ij}=\mathcal{R}(s_{i},s_{j})$.
Figure 1: (Left) Density of the bridge distribution with $\phi=(1+3/\pi^{2})^{-1/2}$ (solid) and normal distribution (dashed), both with unit variance. (Center, Right) Two different bivariate distributions with bridge-distributed marginals with parameter $\phi$ based on the scale mixture of normal and Gaussian copula, both with correlation 0.8.
The requirement that $R$ must have a unit diagonal is crucial, which ensures bridge-distributed marginals through the Theorem 1. By the scale mixture of multivariate normal construction, the multivariate bridge distribution is elliptically symmetric (Fang et al.,, 1990); Figure 1 shows a comparison with a copula-based formulation. Consequently, the correlation structure of the bridge process $u$ aligns with the chosen correlation kernel $\mathcal{R}$.
@@
The Proposition 1 implies that popular choices of spatial kernels retain familiar interpretations, such as the range and smoothness parameters of Matérn kernels. This property is attractive since researchers usually focus on second-order information in modeling spatial dependence (Gelfand and Schliep,, 2016). This is in contrast to the copula-based formulation, where the induced dependence structure of the spatial random effect is obscure due to the complex nonlinear transformations.
@@
Further details on multivariate bridge distributions are in Appendix A.3.
@@ 2.3 Spatial random effect modeling with bridge processes
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid x_{ij},u(s_{i})\}]=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i}),\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (4)
for $i=1,\dots,n$ and $j=1,\dots,N_{i}$. The model (4) induces a marginal logistic regression with coefficient $\phi\beta$, that is, $\mathrm{logit}\{\operatorname*{pr}(Y_{ij}=1\mid x_{ij})\}=\phi x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta$. Thus, the parameter $\phi\in(0,1)$ that determines the variance of random effects can be considered as an attenuation factor, where $\beta$ and $\phi\beta$ carry site-specific and population-averaged interpretations, respectively.
To provide a clearer picture of the induced dependence structure between binary outcomes, we introduce an equivalent representation of (4) based on the thresholding of latent variables.
@@
$Y_{ij}=1(Z_{ij}>0),\quad Z_{ij}=x_{ij}^{\mathrm{\scriptscriptstyle T}}\beta+u(s_{i})+\epsilon_{ij},\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (5)
@@
By multiplying $\phi$ at both sides of the second display of (5), it is evident that $\phi\beta$ carries population-averaged interpretation since $\phi u(s_{i})+\phi\epsilon_{ij}$ follows the standard logistic distribution.
3 Posterior inference strategies
3.1 Partially collapsed Gibbs sampler
Many existing inference methods for spatial logistic regression models, or in general for spatial generalized linear mixed models, are based on Gaussian process spatial random effects. For example, integrated nested Laplace approximation (Rue et al.,, 2009; Lindgren et al.,, 2011) handles a non-Gaussian response variable with Laplace approximation together with latent Gaussian structure to carry out fast approximate inference. When random effects no longer follow Gaussian processes to preserve population-averaged interpretation of coefficients, existing approaches cannot be easily applied; instead, based on Markov chain Monte Carlo, generic sampling algorithms such as Metropolis-Hastings are often adopted (Gory et al.,, 2021). However, such an algorithm is sensitive to tuning and is not appropriate for sampling high-dimensional spatial random effects.
Under a Bayesian framework, we describe how the posterior inference for the proposed model (4) can be efficiently carried out, leveraging normal mixture representation of the bridge distribution. For simplicity, we focus on prior $p(\beta,\phi)=p(\beta)p(\phi)$ with a normal prior for $\beta$, some proper prior on $\phi\in(0,1)$, and a known correlation kernel $\mathcal{R}$, but the extension to inference under a normal scale mixture shrinkage prior or inference on unknown parameter(s) of $\mathcal{R}$ are straightforward by adding additional sampling steps.
[1] Sample $\beta\sim[\beta\mid\omega^{\mathrm{(old)}},\lambda^{\mathrm{(old)}}]$ from multivariate normal, where $u_{1:n}$ is collapsed out.
[2] Jointly sample $(\phi,\lambda,u_{1:n})\sim[\phi,\lambda,u_{1:n}\mid\omega^{\mathrm{(old)}},\beta]$ in two steps:
 [2(i)] Sample $(\phi,\lambda)\sim[\phi,\lambda\mid\omega^{\mathrm{(old)}},\beta]$ using particle marginal Metropolis–Hastings,
 [2(ii)] Sample $u_{1:n}\sim[u_{1:n}\mid\phi,\lambda,\omega^{\mathrm{(old)}},\beta]$ from multivariate normal.
[3] Sample $\omega_{ij}\sim[\omega_{ij}\mid\beta,u_{1:n}]$ from Pólya-Gamma, independently for all $i,j$.
Algorithm 1 One cycle of a partially collapsed Gibbs sampler.
The Algorithm 1 describes a partially collapsed Gibbs sampler (Van Dyk and Park,, 2008), and we outline key strategies involved. Writing $Y=\{Y_{ij}\}$, $X=\{x_{ij}\}$, $u_{1:n}=\{u(s_{1}),\dots,u(s_{n})\}$, augmented data $\omega=\{\omega_{ij}\}$ and $p_{\textsc{pg}}$ as the density of the Pólya-Gamma$(1,0)$ distribution, the complete data model via Pólya-Gamma augmentation (Polson et al.,, 2013) and prior becomes

### v3

@@ 1 Introduction
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid\mathbf{x}_{ij},u(s_{i})\}]=\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i}),\quad u(\cdot)\sim\textrm{mean zero process},$ (1)
@@
where $Y_{ij}$ is the $j$th binary observation at the $i$th spatial location $s_{i}$ defining clusters ($i=1,\dots,n$, $j=1,\dots,N_{i}$), $\mathbf{x}_{ij}\in\mathbb{R}^{p}$ are covariates, and $\{u(s):s\in\mathscr{S}\}$ is a mean zero stochastic process defined on a spatial domain $\mathscr{S}$, with the most common choice being the Gaussian process. This generalized linear mixed effects model formulation (Diggle et al., , 1998) provides a full characterization of the data, naturally accounting for spatial dependence, and allowing for prediction at new locations in a coherent probabilistic framework.
@@
When interpreting the regression coefficient $\bm{\beta}$, the logit link provides a familiar understanding in terms of log odds ratios. However, the inclusion of random effects changes the meaning of $\bm{\beta}$ from a marginal effect on the population to a conditional effect specific to the site. Specifically, when the normal random effect $u$ is integrated out from (1), the induced model for $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid\mathbf{x}_{ij})\}$ is no longer a linear function of $\mathbf{x}_{ij}$ (Zeger et al., , 1988). There exists a rich literature on methods designed for marginally specified models, most notably generalized estimating equations (Liang and Zeger, , 1986), but such approaches cannot provide probabilistic prediction at new locations. This creates difficulty for researchers who wish to maintain a population-averaged interpretation of $\bm{\beta}$ while also accounting for spatial dependence and making predictions. We refer the reader to Neuhaus et al., (1991); Heagerty and Zeger, (2000); Hubbard et al., (2010) for a detailed comparison of marginal and conditional models for clustered binary data.
@@
For logistic models with independent random intercepts, Wang and Louis, (2003) proposed a family of univariate bridge distributions as an alternative to normal random effects. In marginalizing random intercept logistic regression models over the bridge distribution, the resulting model also has a logistic form so that $\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid\mathbf{x}_{ij})\}$, where $u(s_{i})$ is integrated out, is a linear function of $\mathbf{x}_{ij}$ with coefficients proportional to $\bm{\beta}$. Thus, bridge-distributed random effects allow both a conditional and marginal interpretation of regression coefficients, motivating applications in many different contexts (Bandyopadhyay et al., , 2010; Tu et al., , 2011; Asar, , 2021). However, the bridge distribution has not been naturally extended to multivariate settings for correlated random effects.
@@
We propose a new class of marginally interpretable spatial logistic regression models based on a novel spatial bridge process. We identify the normal-scale mixture representation of the bridge distribution and propose its multivariate extension. In contrast to existing copula-based constructions that have bridge-distributed marginals (Lin et al., , 2010; Li et al., , 2011; Parzen et al., , 2011; Boehm et al., , 2013; Swihart et al., , 2014), the bridge process has appealing properties, such as transparent correlation structure and comes with significant computational benefits. We defer all proofs to Supplementary Section A.1.
@@ 2 Marginally interpretable spatial logistic regression models
2.1 Review of bridge distribution and random intercept logistic models
We first review bridge-distributed random intercept logistic regression models without spatial dependence. Wang and Louis, (2003) studied a class of random-intercept distributions with density ${p_{\textsc{br}}}(u;\phi)$ such that, after integrating out the random intercepts, the resulting marginal model remains logistic with its coefficients multiplied by $\phi$. This amounts to finding ${p_{\textsc{br}}}(u;\phi)$ satisfying the integral identity $\int_{-\infty}^{\infty}\mathrm{logit}^{-1}(\eta+u){p_{\textsc{br}}}(u;\phi)\mathrm{d}u=\mathrm{logit}^{-1}(\phi\eta)$ for any $\eta\in\mathbb{R}$, where $\phi$ must satisfy $\phi\in(0,1)$. The corresponding distribution, called the bridge distribution with logit link, has density
${p_{\textsc{br}}}(u;\phi)=\sin(\phi\pi)/[2\pi\{\cosh(\phi u)+\cos(\phi\pi)\}],\quad u\in\mathbb{R}$
with parameter $\phi\in(0,1)$. See Figure 1 (left) for a comparison with the normal distribution. Under the bridge-distributed random intercept logistic model, we have
$\operatorname*{logit}[\operatorname*{pr}\{Y_{ij}=1\mid\mathbf{x}_{ij},u_{i}\}]$ $=\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u_{i},\quad u_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{p_{\textsc{br}}}(u;\phi)$ (2)
$\operatorname*{logit}\{\operatorname*{pr}(Y_{ij}=1\mid\mathbf{x}_{ij})\}$ $=\phi\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}=\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\bm{\beta}^{\textsc{m}}}$ (3)
for $i=1,\dots,n$, $j=1,\dots,N_{i}$, where (2) implies (3) after integrating out $u_{i}$ so that ${\bm{\beta}^{\textsc{m}}}=\phi\bm{\beta}$ has an explicit marginal, population-averaged interpretation, where $\phi$ serve as an attenuation factor (Gory et al., , 2021). Although the marginal probability of $Y_{ij}$ in equation (3) depends only on $\phi\bm{\beta}$, the parameter $\phi$ separately determines the variance of the random intercept. This implies that, after integrating out $u_{i}$, $\phi$ governs the degree of overdispersion in the cluster sums $\sum_{j=1}^{N_{i}}Y_{ij}$ relative to the binomial distribution (Wang and Louis, , 2004), which is formally stated in the following proposition in a simplified setting.
Proposition 1.
Assume that covariates are constant within the cluster, i.e. $\mathbf{x}_{ij}\equiv\mathbf{x}_{i}$ for $j=1,\dots,N_{i}$. Then $\phi=E(\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i},u_{i}))/\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})$, the proportion of the variance of $Y_{ij}$ that is not due to the variability of $u_{i}$, and
${\operatorname*{var}\left(\sum_{j=1}^{N_{i}}Y_{ij}\mid\mathbf{x}_{i}\right)=N_{i}\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})\{1+(N_{i}-1)(1-\phi)\}}$
Thus, as $\phi$ becomes closer to 1, the within-cluster correlation induced by the random intercept disappears, and the model reduces to an ordinary logistic regression model with independent observations, i.e. $\operatorname*{var}(\sum_{j=1}^{N_{i}}Y_{ij}\mid\mathbf{x}_{i})=N_{i}\operatorname*{var}(Y_{ij}\mid\mathbf{x}_{i})$.
2.2 Normal scale mixture representation
For analyzing spatial binary data, we aim to develop a spatial extension of the bridge random intercept model (2) with appealing properties, including a dual interpretation of coefficients. This requires constructing a mean-zero stochastic process for the spatial random effect whose finite-dimensional distributions have bridge-distributed marginals. We first identify a scale mixture of normal representation of the bridge distribution, which forms the basis of our construction for multivariate extensions.
@@ 2.2 Normal scale mixture representation
with the mixing variable $\lambda$ with density ${p_{\textsc{m}}}(\lambda;\phi)$ is equal in distribution to $2\phi^{-2}\sum_{k=1}^{\infty}A_{k}B_{k}/k^{2}$, where $A_{k}$ is an exponential random variable with mean 1, $B_{k}$ is a Bernoulli (binary) random variable mean $1-\phi^{2}$, both independently for $k\in\mathbb{N}$.
Our key contribution is the explicit characterization of the mixing distribution, denoted as $\lambda\sim{p_{\textsc{m}}}(\phi)$; the proof is similar to West, (1987). The mixing distribution has a similar form to the squared Kolmogorov distribution that serves as a normal variance mixing distribution of the logistic distribution (Andrews and Mallows, , 1974), except for the presence of Bernoulli random variables $\{B_{k}\}$. More details of the mixing distribution, such as its density, can be found in Supplementary Section A.3.
@@ 2.3 Bridge processes for logit link
We now introduce a multivariate extension of the univariate bridge distribution, which is uniquely determined by $\phi$ and the choice of a correlation kernel $\mathcal{R}$.
Definition 1 (Bridge processes).
Let $\mathcal{R}:\mathscr{S}\times\mathscr{S}\to[-1,1]$ be a positive semidefinite kernel with $\mathcal{R}(s,s)=1$ for every $s\in\mathscr{S}$. We say $\{u(s)\in\mathbb{R}:s\in\mathscr{S}\}$ is a bridge process with parameter $\phi\in(0,1)$ and correlation kernel $\mathcal{R}$ if every finite collection $\mathbf{u}_{1:n}=\{u(s_{1}),\dots,u(s_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows the multivariate bridge distribution defined as
$\mathbf{u}_{1:n}\,|\,\lambda\sim\mathrm{N}_{n}\left(0,\lambda\mathbf{R}\right),\quad\lambda\sim{p_{\textsc{m}}}(\phi).$ (5)
where $\mathbf{R}$ is a $n\times n$ matrix with $(i,j)$th element $\mathcal{R}(s_{i},s_{j})$. Here $\mathrm{N}_{n}(\bm{\mu},\Sigma)$ denotes $n$-dimensional multivariate normal distribution with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$.
Figure 1: (Left) Density of the bridge distribution with $\phi=(1+3/\pi^{2})^{-1/2}\approx 0.876$ (solid) and normal distribution (dashed), both with unit variance. (Center, Right) Two different bivariate distributions with bridge-distributed marginals with parameter $\phi$ based on the scale mixture of normal and Gaussian copula, both with correlation 0.8.
The requirement that $\mathcal{R}(s,s)=1$ is crucial, which ensures bridge-distributed marginals through Theorem 1. Another key requirement for a stochastic process to be used for spatial prediction is Kolmogorov consistency, meaning that finite-dimensional realizations (5) must be compatible under marginalization; see Banerjee et al., (2014, §3.1). Further details on multivariate bridge distributions, including their densities, are available in Supplementary Section A.3.
Existing multivariate and process extensions of the bridge distribution rely on copula-based formulations (Lin et al., , 2010; Li et al., , 2011; Parzen et al., , 2011; Boehm et al., , 2013; Swihart et al., , 2014). For example, with a Gaussian copula, one can define a process with bridge-distributed marginals through a nonlinear, coordinatewise transformation of a mean-zero, unit-variance Gaussian process $\zeta(\cdot)$ via $\zeta\mapsto F_{B}^{-1}(\Phi(\zeta);\phi)$, where $\Phi$ is a standard normal c.d.f. and $F_{B}^{-1}(\cdot;\phi)$ is the inverse c.d.f. of the bridge distribution. However, compared to the copula-based formulation, the proposed bridge process in Definition 1 through a scale mixture of normals offers several practical benefits.
First, by the scale mixture of normal construction, the realization of the proposed bridge process is elliptically symmetric (Fang et al., , 1990), meaning that the contours of the probability density form an ellipse/ellipsoids; see Figure 1 for a comparison with a copula-based formulation. This elliptical symmetry leads to the following proposition.
@@
The Proposition 2 implies that popular choices of spatial kernels $\mathcal{R}_{\rho}$ with parameter(s) $\rho$ retain familiar interpretations in terms of correlation between spatial random effects $u(\cdot)$, such as the range and smoothness parameters of Matérn kernels. This is in contrast to the copula-based formulation, where the induced dependence structure via a latent process is obscure due to the complex nonlinear transformations involved, i.e. $\operatorname*{corr}\{F_{B}^{-1}(\Phi(\zeta(s));\phi),F_{B}^{-1}(\Phi(\zeta(s^{\prime}));\phi)\}\neq\operatorname*{corr}\{\zeta(s),\zeta(s^{\prime})\}$ for a latent Gaussian process $\zeta(\cdot)$.
@@
Next, because the bridge process specifies a transparent correlation structure for the spatial random effect, it can be incorporated naturally into the standard model-based geostatistics framework, which consists of specifying a data model, a process model, and a parameter model (Berliner, , 1996; Wikle et al., , 2019). In this framework, the process model introduces dependence structure on the linear predictor scale, typically focusing on second-order information to characterize spatial dependence (Gelfand and Schliep, , 2016). For binary data with a logit link, Diggle and Giorgi, (2016) refer to this formulation as the standard geostatistical model for prevalence data. Finally, the normal scale mixture representation of the bridge process provides significant benefits for posterior computation together with Pólya-Gamma augmentation (Polson et al., , 2013), which will be elaborated in detail in Section 3.2.
@@ 2.4 Spatial random effect modeling with bridge processes
$\mathrm{logit}[\operatorname*{pr}\{Y_{ij}=1\mid\mathbf{x}_{ij},u(s_{i})\}]=\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i}),\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (6)
for $i=1,\dots,n$ and $j=1,\dots,N_{i}$. As in the random intercept model, the proposed model (6) induces a marginal logistic regression model (3), where $\bm{\beta}$ and $\phi\bm{\beta}$ carry site-specific and population-averaged interpretations, respectively. To provide a clearer picture of the induced dependence structure between binary outcomes, we introduce an equivalent representation of (6) based on the thresholding of latent variables.
@@
$Y_{ij}=1(Z_{ij}>0),\quad Z_{ij}=\mathbf{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}+u(s_{i})+\epsilon_{ij},\quad u(\cdot)\sim\mathrm{bridge\,process}(\phi,\mathcal{R}),$ (7)
@@
By multiplying $\phi$ at both sides of the second display of (7), it is evident that $\phi\bm{\beta}$ carries population-averaged interpretation since $\phi u(s_{i})+\phi\epsilon_{ij}$ follows the standard logistic distribution.
3 Posterior inference
3.1 Empirical Bayes approach for attenuation factor
The proposed marginally interpretable spatial logistic model (6) has parameters $(\bm{\beta},\phi,\rho)$, where $\rho$ denotes a correlation kernel parameter. The complete data likelihood, including all latent variables, is given by